# BeamDojo: Learning Agile Humanoid Locomotion on Sparse Footholds

### v2

@@ Abstract
Traversing risky terrains with sparse footholds poses a significant challenge for humanoid robots, requiring precise foot placements and stable locomotion. Existing approaches designed for quadrupedal robots often fail to generalize to humanoid robots due to differences in foot geometry and unstable morphology, while learning-based approaches for humanoid locomotion still face great challenges on complex terrains due to sparse foothold reward signals and inefficient learning processes. To address these challenges, we introduce BeamDojo, a reinforcement learning (RL) framework designed for enabling agile humanoid locomotion on sparse footholds. BeamDojo begins by introducing a sampling-based foothold reward tailored for polygonal feet, along with a double critic to balancing the learning process between dense locomotion rewards and sparse foothold rewards. To encourage sufficient trail-and-error exploration, BeamDojo incorporates a two-stage RL approach: the first stage relaxes the terrain dynamics by training the humanoid on flat terrain while providing it with task terrain perceptive observations, and the second stage fine-tunes the policy on the actual task terrain. Moreover, we implement a onboard LiDAR-based elevation map to enable real-world deployment. Extensive simulation and real-world experiments demonstrate that BeamDojo achieves efficient learning in simulation and enables agile locomotion with precise foot placement on sparse footholds in the real world, maintaining a high success rate even under significant external disturbances.
@@ I Introduction
Existing works have effectively addressed this complex task for quadrupedal robots [14, 15, 26, 34, 53, 58, 59, 61, 62]. However, these methods encounter great challenges when applied to humanoid robots, primarily due to a key difference in foot geometry. Although the foot of most quadrupedal robots and some simplified bipedal robots [8, 29] can be modeled as a point, the foot of humanoid robots is often represented as a polygon [5, 16, 23, 47]. For traditional model-based methods, this requires additional half-space constraints defined by linear inequalities, which impose a significant computational burden for online planning [7, 16, 37, 47]. In the case of reinforcement learning (RL) methods, foothold rewards designed for point-shaped feet are not suitable for evaluating foot placement of polygon-shaped feet [58]. Hybrid methods, which combine RL with model-based controllers, face similar challenges in online planning for humanoid feet [14, 26, 53]. Furthermore, the higher degrees of freedom and the inherently unstable morphology of humanoid robots make it even more difficult to achieve agile and stable locomotion over risky terrains.
@@ II-A Locomotion on Sparse Footholds
Walking on sparse footholds has been a long-standing application of perceptive legged locomotion. Existing works often employs model-based hierarchical controllers, which decompose this complex task into separate stages of perception, planning, and control [15, 16, 25, 39, 40, 51]. However, model-based controllers react sensitively to violation of model assumptions, which hinders applications in real-world scenarios. Recent studies have explored combining RL with model-based controllers, such as using RL to generate trajectories that are then tracked by model-based controllers [14, 59, 53], or employing RL policies to track trajectories generated by model-based planners [26]. While demonstrating remarkable performance, these decoupled architectures can constrain the adaptability and coordination of each module.
@@ IV-C Learning Terrain-Aware Locomotion via Two-Stage RL
To address the early termination problem in complex terrain dynamics and encourage full trial-and-error exploration, we adopt a two-stage reinforcement learning (RL) approach for terrain-aware locomotion in simulation, inspired by [63, 64]. As illustrated in Fig. 3, in the first stage, termed the “soft terrain dynamics constraints” phase, the humanoid robot is trained on flat terrain while being provided with a corresponding height map of the true task terrains (e.g., stepping stones). This setup encourages broad exploration without the risk of early termination from missteps. Missteps are penalized but do not lead to termination, allowing the humanoid robot to develop foundational skills for terrain-aware locomotion. In the second stage, termed the “hard terrain dynamics constraints” phase, we continue training the humanoid on the real terrains in simulation, where missteps result in termination. This stage fine-tunes the robot’s ability to step on challenging terrains accurately.
@@ V-A Experimental Setup
  1. 1.
    PIM [32]: This one-stage method is designed for humanoid locomotion tasks, such as walking up stairs and traversing uneven terrains. We additional add our foothold reward $r_{\text{foothold}}$ to encourage the humanoid to step accurately on the foothold areas.
  2. 2.
@@
The training and simulation environments are implemented in IsaacGym [35]. To ensure fairness, we adapt all methods to two stages. For stage 1, we train the humanoid on the Stones Everywhere with curriculum learning. In this stage, our method and baseline 4 use soft terrain dynamics constraints, while all other baselines use hard terrain dynamics constraints. For stage 2, we conduct fine-tuning on the Stepping Stones and Balancing Beams terrains with curriculum learning.
@@ V-B1 Quantitative results
  • Despite the our method not being explicitly trained on Stepping Beams and Gaps, it demonstrates impressive zero-shot generalization capabilities on this terrain.
@@ V-B2 Detailed Ablation Analysis
We conduct additional ablation studies by comparing BeamDojo with baselines 23, and 4.
@@ V-C1 Result
As demonstrated in Fig.1, our framework achieves zero-shot transfer, successfully generalizing to real-world dynamics. To showcase the effect of height map domain randomization (introduced in Section IV-D3) in sim-to-real transfer, we compare our proposed method with an ablation that excludes height map randomization (denoted as ”ours w/o HR”). We conduct five trials on each terrain and report the success and traversal rates in Fig. 8, with following conclusions:
@@ V-C2 Agility Test
To assess the agility of our method, we provide the humanoid robot with three commanded longitudinal velocities, $\mathbf{v}_{x}^{c}$: 0.5, 0.75, and 1.0 m/s, and check the tracking ability. Each test was conducted over three trials, and the results are reported in Table IV. The results show minimal tracking error, even at the highest command velocity of 1.0 m/s, where the robot achieved an average speed of 0.88 m/s, which demonstrate the agility of our policy.
@@ V-D3 Design of Commands
V-D4 Failure Cases and Limitations
@@ Acknowledgments
We thank the authors of PIM [32] for their kind help with the deployment of the elevation map. We thank Unitree Robotics for their help with the G1 robot.

### v3

@@ Abstract
Traversing risky terrains with sparse footholds poses a significant challenge for humanoid robots, requiring precise foot placements and stable locomotion. Existing learning-based approaches often struggle on such complex terrains due to sparse foothold rewards and inefficient learning processes. To address these challenges, we introduce BeamDojo, a reinforcement learning (RL) framework designed for enabling agile humanoid locomotion on sparse footholds. BeamDojo begins by introducing a sampling-based foothold reward tailored for polygonal feet, along with a double critic to balancing the learning process between dense locomotion rewards and sparse foothold rewards. To encourage sufficient trial-and-error exploration, BeamDojo incorporates a two-stage RL approach: the first stage relaxes the terrain dynamics by training the humanoid on flat terrain while providing it with task-terrain perceptive observations, and the second stage fine-tunes the policy on the actual task terrain. Moreover, we implement a onboard LiDAR-based elevation map to enable real-world deployment. Extensive simulation and real-world experiments demonstrate that BeamDojo achieves efficient learning in simulation and enables agile locomotion with precise foot placement on sparse footholds in the real world, maintaining a high success rate even under significant external disturbances.
@@ I Introduction
Existing works have effectively addressed this complex task for quadrupedal robots [15, 16, 27, 35, 55, 60, 61, 63, 64]. However, these methods encounter great challenges when applied to humanoid robots, primarily due to a key difference in foot geometry. Although the foot of most quadrupedal robots and some simplified bipedal robots [9, 30] can be modeled as a point, the foot of humanoid robots is often represented as a polygon [5, 17, 24, 49]. For traditional model-based methods, this requires additional half-space constraints defined by linear inequalities, which impose a significant computational burden for online planning [7, 8, 17, 38, 49]. In the case of reinforcement learning (RL) methods, foothold rewards designed for point-shaped feet are not suitable for evaluating foot placement of polygon-shaped feet [60]. Hybrid methods, which combine RL with model-based controllers, face similar challenges in online planning for humanoid feet [15, 27, 55]. Furthermore, the higher degrees of freedom and the inherently unstable morphology of humanoid robots make it even more difficult to achieve agile and stable locomotion over risky terrains.
@@ II-A Locomotion on Sparse Footholds
Walking on sparse footholds has been a long-standing application of perceptive legged locomotion. Existing works often employs model-based hierarchical controllers, which decompose this complex task into separate stages of perception, planning, and control [16, 17, 26, 40, 41, 43, 53]. However, model-based controllers react sensitively to violation of model assumptions, which hinders applications in real-world scenarios. Recent studies have explored combining RL with model-based controllers, such as using RL to generate trajectories that are then tracked by model-based controllers [15, 61, 55], or employing RL policies to track trajectories generated by model-based planners [27]. While demonstrating remarkable performance, these decoupled architectures can constrain the adaptability and coordination of each module.
@@ IV-C Learning Terrain-Aware Locomotion via Two-Stage RL
To address the early termination problem in complex terrain dynamics and encourage full trial-and-error exploration, we design a novel two-stage reinforcement learning (RL) approach for terrain-aware locomotion in simulation. As illustrated in Fig. 3, in the first stage, termed the “soft terrain dynamics constraints” phase, the humanoid robot is trained on flat terrain while being provided with a corresponding height map of the true task terrains (e.g., stepping stones). This setup encourages broad exploration without the risk of early termination from missteps. Missteps are penalized but do not lead to termination, allowing the humanoid robot to develop foundational skills for terrain-aware locomotion. In the second stage, termed the “hard terrain dynamics constraints” phase, we continue training the humanoid on the real terrains in simulation, where missteps result in termination. This stage fine-tunes the robot’s ability to step on challenging terrains accurately.
@@ V-A Experimental Setup
  1. BL 1)
    PIM [34]: This one-stage method is designed for humanoid locomotion tasks, such as walking up stairs and traversing uneven terrains. We additionally add our foothold reward $r_{\text{foothold}}$ to encourage the humanoid to step accurately on the foothold areas.
  2. BL 2)
@@ 4. BL 4)
The training and simulation environments are implemented in IsaacGym [36]. To ensure fairness, we adapt all methods to two stages. For stage 1, we train the humanoid on the Stones Everywhere with curriculum learning. In this stage, our method and BL 4) use soft terrain dynamics constraints, while all other baselines use hard terrain dynamics constraints. For stage 2, we conduct fine-tuning on the Stepping Stones and Balancing Beams terrains with curriculum learning.
@@ V-B1 Quantitative results
  • Despite the our method not being explicitly trained on Stepping Beams and Gaps, it demonstrates impressive zero-shot generalization capabilities on these terrains.
@@ V-B2 Detailed Ablation Analysis
We conduct additional ablation studies by comparing BeamDojo with BL 2), BL 3), and BL 4).
@@ V-C1 Result
As demonstrated in Fig.1, our framework achieves zero-shot transfer, successfully generalizing to real-world dynamics. To showcase the effect of height map domain randomization (introduced in Section IV-D3) in sim-to-real transfer, we compare our proposed method with an ablation that excludes height map randomization (denoted as “ours w/o HR”). We conduct five trials on each terrain and report the success and traversal rates in Fig. 8, with following conclusions:
@@ V-C2 Agility Test
$1.25$ $2.91$ (±$\pm 0.63$) $0.96$ (±$\pm 0.03$) $22.74$ (±$\pm 5.32$)
$1.5$ $2.69$ (±$\pm 0.42$) $1.04$ (±$\pm 0.05$) $30.68$ (±$\pm 6.17$)
To assess the agility of our method, we provide the humanoid robot with five commanded longitudinal velocities, $\mathbf{v}_{x}^{c}$: 0.5, 0.75, 1.0, 1.25, and 1.5 m/s, and check the tracking ability. Each test was conducted over three trials, and the results are reported in Table IV. The results show minimal tracking error up to the highest training command velocity of 1.0 m/s, where the robot achieves an average speed of 0.88 m/s, demonstrating the agility of our policy. However, performance degrades significantly beyond 1.25 m/s, as maintaining such speeds becomes increasingly difficult on these highly challenging terrains.
@@ V-D3 Design of Commands
V-D4 Generalization to Non-Flat Terrains
Figure 10: Generalization Test on Non-Flat Terrains. We conduct real-world experiments on (a) stairs with a width of 25 cm and a height of 15 cm, and (b) slopes with a 15-degree incline.
We observe that BeamDojo also generalizes well to non-flat terrains such as stairs and slopes. The main adaptation involves calculating the base height reward relative to the foot height rather than the ground height on these uneven surfaces. Additionally, Stage,1 pre-training becomes unnecessary for stairs and slopes, as the footholds are no longer sparse. We validate our approach through hardware experiments on stairs and slopes, as shown in Fig. 10. The results demonstrate that BeamDojo enables the robot to successfully traverse stairs and slopes with success rates of 8/10 and 10/10, respectively.
V-D5 Failure Cases
Figure 11: Failure Case Analysis. We evaluate the success rate on varying (a) stone sizes, and (b) step distances.
To investigate the framework’s performance limitations, we evaluate its performance across varying stone sizes and step distances, as shown in Fig. 11. We compare policies trained with different terrain parameters, with the smallest stone size (20 cm, 15 cm and 10 cm) or largest step distance (45 cm, 50 cm and 55 cm). The results indicate that while tougher training enhances adaptability, performance still drops sharply on 10 cm stones (approximately half the foot length) and 55 cm steps (roughly equal to the leg length), even under the most challenging training settings. In these cases, the difficulty shifts toward maintaining balance on very small footholds and executing larger strides—challenges that the current reward function does not adequately address.
V-D6 Limitations
@@ Acknowledgments
This work is funded in part by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory. We thank the authors of PIM [34] for their kind help with the deployment of the elevation map. We thank Unitree Robotics for their help with the Unitree G1 humanoid robot. We thank RSS reviewers for their careful and insightful feedback, which helped improve the quality of this work.